# Modelling and calibration of pair-rule protein patterns in Drosophila embryo: From Even-skipped and Fushi-tarazu to Wingless expression networks

### v3

@@ Abstract
We modelled and calibrated the distributions of the seven-stripe patterns of Even-skipped (Eve) and Fushi-tarazu (Ftz) pair-rule proteins along the anteroposterior axis of the Drosphila embryo, established during early development. We have identified the putative repressive combinations for five Eve enhancers, and we have explored the relationship between Eve and Ftz for complementary patterns. The regulators of Eve and Ftz are stripe-specific DNA enhancers with embryo position-dependent activation rates and regulated by the gap family of proteins. We achieved remarkable data matching of the Eve stripe pattern, and the calibrated model reproduces gap gene mutation experiments. Extended work inferring the Wingless (Wg) fourteen stripe pattern from Eve and Ftz enhancers have been proposed, clarifying the hierarchical structure of Drosphila’s genetic expression network during early development.
@@ A.1 Transcriptional regulation model for stripes 1 and 2 of Eve, and dual stripe 1+5 of Ftz
$F(\cdot)=\left(\frac{\overline{\text{G}}(0)}{\alpha_{0}}+\frac{\overline{\text{Kr}}(t)-\overline{\text{Kr}}(0)}{\alpha_{2}}+\frac{\overline{\text{Kni}}(t)-\overline{\text{Kni}}(0)}{\alpha_{3}}+\frac{\overline{\text{Gt}}(t)-\overline{\text{Gt}}(0)}{\alpha_{4}}+\frac{\overline{\text{Tll}}(t)-\overline{\text{Tll}}(0)}{\alpha_{6}}+\frac{\overline{\text{Hkb}}(t)-\overline{\text{Hkb}}(0)}{\alpha_{7}}\right),$
@@ A.2 Transcriptional regulation model for dual stripes 3+7 of Eve, and dual stripe 2+7 and 3+6 of Ftz
$F(\cdot)=\left(\frac{\overline{\text{G}}(0)}{\alpha_{0}}+\frac{\overline{\text{Kni}}(t)-\overline{\text{Kni}}(0)}{\alpha_{3}}+\frac{\overline{\text{Hb}}(t)-\overline{\text{Hb}}(0)}{\alpha_{5}}+\frac{\overline{\text{Tll}}(t)-\overline{\text{Tll}}(0)}{\alpha_{6}}+\frac{\overline{\text{Hkb}}(t)-\overline{\text{Hkb}}(0)}{\alpha_{7}}\right).$
@@ A.3 Transcriptional regulation model for dual stripes 4+6 of Eve
$F(\cdot)=\left(\frac{\overline{\text{G}}_{e}(0)}{\alpha_{0}}+\frac{\overline{\text{Kni}}(t)-\overline{\text{Kni}}(0)}{\alpha_{3}}+\frac{\overline{\text{Gt}}(t)-\overline{\text{Gt}}(0)}{\alpha_{4}}+\frac{\overline{\text{Hb}}(t)-\overline{\text{Hb}}(0)}{\alpha_{5}}+\frac{\overline{\text{Tll}}(t)-\overline{\text{Tll}}(0)}{\alpha_{6}}+\frac{\overline{\text{Hkb}}(t)-\overline{\text{Hkb}}(0)}{\alpha_{7}}\right).$
@@ A.4 Transcriptional regulation model for stripe 5 of Eve
$F(\cdot)=\left(\frac{\overline{\text{G}}_{e}(0)}{\alpha_{0}}+\frac{\overline{\text{Kr}}(t)-\overline{\text{Kr}}(0)}{\alpha_{2}}+\frac{\overline{\text{Gt}}(t)-\overline{\text{Gt}}(0)}{\alpha_{4}}+\frac{\overline{\text{Hb}}(t)-\overline{\text{Hb}}(0)}{\alpha_{5}}+\frac{\overline{\text{Tll}}(t)-\overline{\text{Tll}}(0)}{\alpha_{6}}+\frac{\overline{\text{Hkb}}(t)-\overline{\text{Hkb}}(0)}{\alpha_{7}}\right).$
@@ Appendix B Alternative pattern formation: eve-ftz epistasis
$\begin{array}[]{lcl}\overline{\text{Eve}}^{\prime}(t)&=&-\overline{\text{Eve}}(t)\left(k_{3}+\dfrac{\overline{\text{G}}_{f}(t)k_{5}}{\alpha_{1}}\right)+\dfrac{\overline{\text{G}}_{e}(t)k_{1}\alpha_{2}}{\alpha_{0}}+k_{-5}\alpha_{2}F_{1}(\cdot)\\[5.0pt] \overline{\text{Ftz}}^{\prime}(t)&=&-\overline{\text{Ftz}}(t)\left(k_{4}+\dfrac{\overline{\text{G}}_{e}(t)k_{6}}{\alpha_{0}}\right)+\alpha_{3}\left(k_{-6}\text{G}_{R2}(t)+\dfrac{\overline{G}_{f}(t)k_{2}}{\alpha_{1}}\right)\\[5.0pt] \overline{\text{G}}_{e}^{\prime}(t)&=&-\overline{\text{G}}_{e}(t)\left(\dfrac{\overline{\text{Ftz}}(t)k_{6}}{\alpha_{3}}+\dfrac{\overline{\text{TR}}(t)k_{7}}{\alpha_{4}}\right)+k_{-6}\alpha_{0}\text{G}_{R2}(t)-k_{-7}\alpha_{0}F_{2}(\cdot)\\[5.0pt] \overline{\text{G}}_{f}^{\prime}(t)&=&-\overline{\text{G}}_{f}(t)\left(\dfrac{\overline{\text{Eve}}(t)k_{5}}{\alpha_{2}}+\dfrac{\overline{\text{TR}}(t)k_{8}}{\alpha_{4}}\right)+k_{-5}\alpha_{1}F_{1}(\cdot)+k_{-8}\alpha_{1}F_{3}(\cdot)\\[5.0pt] \overline{\text{TR}}^{\prime}(t)&=&-\overline{\text{TR}}(t)\left(\dfrac{\overline{\text{G}}_{e}(t)k_{7}}{\alpha_{0}}+\dfrac{\overline{\text{G}}_{f}(t)k_{8}}{\alpha_{1}}\right)-k_{-7}\alpha_{4}F_{2}(\cdot)+k_{-8}\alpha_{4}F_{3}(\cdot)\\[5.0pt] \text{G}_{R2}^{\prime}(t)&=&-\text{G}_{R2}(t)k_{-6}+\dfrac{\overline{\text{Ftz}}(t)\overline{G}_{e}(t)k_{6}}{\alpha_{0}\alpha_{3}},\end{array}$
@@
$\begin{array}[]{lcl}F_{1}(\cdot)=\left(\text{G}_{R1}(0)+\text{G}_{R2}(0)-\text{G}_{R2}(t)+\dfrac{\overline{\text{G}}_{e}(0)-\overline{\text{G}}_{e}(t)}{\alpha_{0}}+\dfrac{\overline{\text{G}}_{f}(0)-\overline{\text{G}}_{f}(t)}{\alpha_{1}}+\dfrac{\overline{\text{TR}}(t)-\overline{\text{TR}}(0)}{\alpha_{4}}\right)\\[5.0pt] F_{2}(\cdot)=\left(\text{G}_{R2}(0)-\text{G}_{R3}(0)-\text{G}_{R2}(t)+\dfrac{\overline{\text{G}}_{f}(0)-\overline{\text{G}}_{f}(t)}{\alpha_{1}}+\dfrac{\overline{\text{TR}}(t)-\overline{\text{TR}}(0)}{\alpha_{4}}\right)\\[5.0pt] F_{3}(\cdot)=\left(\text{G}_{R2}(0)+\text{G}_{R4}(0)-\text{G}_{R2}(t)+\dfrac{\overline{\text{G}}_{f}(0)-\overline{\text{G}}_{f}(t)}{\alpha_{1}}\right).\end{array}$

### v4

@@ Abstract
We modelled and calibrated the distributions of the seven-stripe patterns of Even-skipped (Eve) and Fushi-tarazu (Ftz) pair-rule proteins along the anteroposterior axis of the Drosphila embryo, established during early development. We have identified the putative repressive combinations for five Eve enhancers, and we have explored the relationship between Eve and Ftz for complementary patterns. The regulators of Eve and Ftz are stripe-specific DNA enhancers with embryo position-dependent activation rates and are regulated by the gap family of proteins. We achieved remarkable data matching of the Eve stripe pattern, and the calibrated model reproduces gap gene mutation experiments. Extended work inferring the Wingless (Wg) fourteen stripe pattern from Eve and Ftz enhancers have been proposed, clarifying the hierarchical structure of Drosphila’s genetic expression network during early development.
@@ A.1 Transcriptional regulation model for stripes 1 and 2 of Eve, and dual stripe 1+5 of Ftz
$\begin{array}[]{ll}F(\cdot)=&\left(\frac{\overline{\text{G}}(0)}{\alpha_{0}}+\frac{\overline{\text{Kr}}(t)-\overline{\text{Kr}}(0)}{\alpha_{2}}+\frac{\overline{\text{Kni}}(t)-\overline{\text{Kni}}(0)}{\alpha_{3}}+\frac{\overline{\text{Gt}}(t)-\overline{\text{Gt}}(0)}{\alpha_{4}}\right.\\ &\left.+\frac{\overline{\text{Tll}}(t)-\overline{\text{Tll}}(0)}{\alpha_{6}}+\frac{\overline{\text{Hkb}}(t)-\overline{\text{Hkb}}(0)}{\alpha_{7}}\right),\end{array}$
@@ A.2 Transcriptional regulation model for dual stripes 3+7 of Eve, and dual stripe 2+7 and 3+6 of Ftz
$\begin{array}[]{ll}F(\cdot)=&\left(\frac{\overline{\text{G}}(0)}{\alpha_{0}}+\frac{\overline{\text{Kni}}(t)-\overline{\text{Kni}}(0)}{\alpha_{3}}+\frac{\overline{\text{Hb}}(t)-\overline{\text{Hb}}(0)}{\alpha_{5}}+\frac{\overline{\text{Tll}}(t)-\overline{\text{Tll}}(0)}{\alpha_{6}}\right.\\ &\left.+\frac{\overline{\text{Hkb}}(t)-\overline{\text{Hkb}}(0)}{\alpha_{7}}\right).\end{array}$
@@ A.3 Transcriptional regulation model for dual stripes 4+6 of Eve
$\begin{array}[]{ll}F(\cdot)=&\left(\frac{\overline{\text{G}}_{e}(0)}{\alpha_{0}}+\frac{\overline{\text{Kni}}(t)-\overline{\text{Kni}}(0)}{\alpha_{3}}+\frac{\overline{\text{Gt}}(t)-\overline{\text{Gt}}(0)}{\alpha_{4}}+\frac{\overline{\text{Hb}}(t)-\overline{\text{Hb}}(0)}{\alpha_{5}}\right.\\ &\left.+\frac{\overline{\text{Tll}}(t)-\overline{\text{Tll}}(0)}{\alpha_{6}}+\frac{\overline{\text{Hkb}}(t)-\overline{\text{Hkb}}(0)}{\alpha_{7}}\right).\end{array}$
@@ A.4 Transcriptional regulation model for stripe 5 of Eve
$\begin{array}[]{ll}F(\cdot)=&\left(\frac{\overline{\text{G}}_{e}(0)}{\alpha_{0}}+\frac{\overline{\text{Kr}}(t)-\overline{\text{Kr}}(0)}{\alpha_{2}}+\frac{\overline{\text{Gt}}(t)-\overline{\text{Gt}}(0)}{\alpha_{4}}+\frac{\overline{\text{Hb}}(t)-\overline{\text{Hb}}(0)}{\alpha_{5}}\right.\\ &\left.+\frac{\overline{\text{Tll}}(t)-\overline{\text{Tll}}(0)}{\alpha_{6}}+\frac{\overline{\text{Hkb}}(t)-\overline{\text{Hkb}}(0)}{\alpha_{7}}\right).\end{array}$
@@ Appendix B Alternative pattern formation: eve-ftz epistasis
$\begin{array}[]{lcl}\overline{\text{Eve}}^{\prime}(t)&=&-\overline{\text{Eve}}(t)\left(k_{3}+\frac{\overline{\text{G}}_{f}(t)k_{5}}{\alpha_{1}}\right)+\frac{\overline{\text{G}}_{e}(t)k_{1}\alpha_{2}}{\alpha_{0}}+k_{-5}\alpha_{2}F_{1}(\cdot)\\[5.0pt] \overline{\text{Ftz}}^{\prime}(t)&=&-\overline{\text{Ftz}}(t)\left(k_{4}+\frac{\overline{\text{G}}_{e}(t)k_{6}}{\alpha_{0}}\right)+\alpha_{3}\left(k_{-6}\text{G}_{R2}(t)+\frac{\overline{G}_{f}(t)k_{2}}{\alpha_{1}}\right)\\[5.0pt] \overline{\text{G}}_{e}^{\prime}(t)&=&-\overline{\text{G}}_{e}(t)\left(\frac{\overline{\text{Ftz}}(t)k_{6}}{\alpha_{3}}+\frac{\overline{\text{TR}}(t)k_{7}}{\alpha_{4}}\right)+k_{-6}\alpha_{0}\text{G}_{R2}(t)-k_{-7}\alpha_{0}F_{2}(\cdot)\\[5.0pt] \overline{\text{G}}_{f}^{\prime}(t)&=&-\overline{\text{G}}_{f}(t)\left(\frac{\overline{\text{Eve}}(t)k_{5}}{\alpha_{2}}+\frac{\overline{\text{TR}}(t)k_{8}}{\alpha_{4}}\right)+k_{-5}\alpha_{1}F_{1}(\cdot)+k_{-8}\alpha_{1}F_{3}(\cdot)\\[5.0pt] \overline{\text{TR}}^{\prime}(t)&=&-\overline{\text{TR}}(t)\left(\frac{\overline{\text{G}}_{e}(t)k_{7}}{\alpha_{0}}+\frac{\overline{\text{G}}_{f}(t)k_{8}}{\alpha_{1}}\right)-k_{-7}\alpha_{4}F_{2}(\cdot)+k_{-8}\alpha_{4}F_{3}(\cdot)\\[5.0pt] \text{G}_{R2}^{\prime}(t)&=&-\text{G}_{R2}(t)k_{-6}+\frac{\overline{\text{Ftz}}(t)\overline{G}_{e}(t)k_{6}}{\alpha_{0}\alpha_{3}},\end{array}$
@@
$\begin{array}[]{lcl}F_{1}(\cdot)&=&\left(\text{G}_{R1}(0)+\text{G}_{R2}(0)-\text{G}_{R2}(t)+\frac{\overline{\text{G}}_{e}(0)-\overline{\text{G}}_{e}(t)}{\alpha_{0}}+\frac{\overline{\text{G}}_{f}(0)-\overline{\text{G}}_{f}(t)}{\alpha_{1}}\right.\\ &&+\left.\frac{\overline{\text{TR}}(t)-\overline{\text{TR}}(0)}{\alpha_{4}}\right)\\[5.0pt] F_{2}(\cdot)&=&\left(\text{G}_{R2}(0)-\text{G}_{R3}(0)-\text{G}_{R2}(t)+\frac{\overline{\text{G}}_{f}(0)-\overline{\text{G}}_{f}(t)}{\alpha_{1}}+\frac{\overline{\text{TR}}(t)-\overline{\text{TR}}(0)}{\alpha_{4}}\right)\\[5.0pt] F_{3}(\cdot)&=&\left(\text{G}_{R2}(0)+\text{G}_{R4}(0)-\text{G}_{R2}(t)+\frac{\overline{\text{G}}_{f}(0)-\overline{\text{G}}_{f}(t)}{\alpha_{1}}\right).\end{array}$